### v1

Abstract
Self-supervised pretraining from static structures of drug-like compounds and proteins enable powerful learned feature representations. Learned features demonstrate state of the art performance on a range of predictive tasks including molecular properties, structure generation, and protein-ligand interactions. The majority of approaches are limited by their use of static structures and it remains an open question, how best to use atomistic molecular dynamics (MD) simulations to develop more generalized models to improve prediction accuracy for novel molecular structures. We present SURrogate mmGBSA (SurGBSA) as a new modeling approach for MD-based representation learning, which learns a surrogate function of the Molecular Mechanics Generalized Born Surface Area (MMGBSA). We show for the first time the benefits of physics-informed pre-training to train a surrogate MMGBSA model on a collection of over 1.4 million 3D trajectories collected from MD simulations of the CASF-2016 benchmark. SurGBSA demonstrates a dramatic 6,497 speedup versus a traditional physics-based single-point MMGBSA calculation while nearly matching single-point MMGBSA accuracy on the challenging pose ranking problem for identification of the correct top pose ( difference). Our work advances the development of molecular foundation models by showing model improvements when training on MD simulations. Models, code and training data are made publicly available.
keywords:
Deep Learning, Representation Learning, Molecular Dynamics, Biophysics, Drug Discovery1 Introduction
Artificial Intelligence (AI) has made a significant impact on computer based drug-design [1, 2, 3, 4, 5]. New tools enable the prediction of structure from sequence [1, 2, 3, 6, 7], binding poses [8, 9], and binding affinities [10, 11, 12, 13, 14]. These advances are driven in part by the development of sample-efficient inductive biases [15, 16], allowing invariance to arbitrary rotation and translation in Euclidean space, coupled with momentum towards the development of increasingly large foundation models [17]. In recent years, foundation models, which are often trained using self-supervision, have enabled significant advances across nearly every application domain of AI and machine learning [17]. Methods that process text and graphs have been further adapted to develop biological foundation models trained using publicly available sequence data for ligands and proteins [18, 19, 20, 21]. In contrast to NLP and Computer Vision, datasets for drug discovery are comparatively limited in the number of training samples, posing unique challenges to foundation model development. Acquiring additional training data remains cost prohibitive on the scales typically observed for foundation model development [17].
Atomistic Molecular Dynamics [22, 23] (MD) has traditionally served as a compliment to physical experiments that use physics theory to sample the dynamics of the protein-ligand co-complex which can then be used to estimate binding free energies in order to prioritize compounds for experimental validation. The trajectories generated by MD contain dynamical aspects of the binding interaction that static representations do not fully capture. However, despite extensive research on MD acceleration, MD simulations continue to require significant computational resources [24], especially when considering the approximately 20,000 protein-ligand co-complex structures currently in the PDB [25], and with generative AI poised to play a significant role in increasing the amount of co-complex data [3]. Consequently, computational effort should be allocated based on the confidence of predicted binding interactions, typically prioritizing the best-scoring poses from molecular docking calculations. Furthermore, the required simulation time varies considerably across different free-energy methods, with some approaches requiring multiple MD trajectories under specific parameter sets [26]. Molecular Mechanics Generalized Borne Surface Area (MMGBSA) and Molecular Mechanics Poission Boltz Surface Area (MMPBSA) serve as alternatives to the more expensive free-energy estimation methods and are commonly used in conjunction with molecular docking in virtual screening to prioritize co-complex structures for more expensive methods. Typically, the MMGBSA and MMPBSA re-scoring methods only require a single MD simulation of the co-complex structure to make their free-energy estimation as the ensemble average over all of the trajectory frames. These methods require generally less simulation time compared to the more expensive Alchemical Free Energy methods. The MMGBSA method makes approximations that increase the efficiency of the calculation considerably [26] and so we based our study on this method specifically. As an additional step for efficiency, “single-point” MMGBSA (and MMPBSA) calculations can be done using a single 3D structure as input, providing a method to re-rank docking poses prior to running more expensive methods [26].
Recent research in molecular representation learning suggests promising results when integrating MD simulations for model development [20, 19, 27, 28, 29, 30, 31, 7, 32]. Libouban et al. [32] investigated training convolution and recurrent architectures on MD simulation frames to augment the PDBBind dataset. The authors show an improvement for predicting binding affinities for the PDBbind v.2016 core set, their MDbind test set, and the FEP dataset. This study did not consider equivariant architectures so it is not clear how these architectures would perform in this context. The earliest work to develop a model trained on MMPBSA energetics was demonstrated by Korlepara et al. [33, 34] based on 4ns MD simulations for 14,500 structures from the PDB [25]. However, the authors of PLAS20k only provide the ensemble-average of the MMPBSA calculation for each PDB co-complex and the models are trained using crystal structures. While the study provides insight into the general problem of learning the MMPBSA energetics, it also doesn’t explicitly model the MD simulation data itself, making it unclear how a model trained using a re-scoring method for label generation, would perform when tasked with re-scoring novel structures in practice. The Misato dataset [29] aggregates MD simulations and quantum mechanics (QM) properties for 19,443 protein-ligand structures from PDBbind [35] while also providing MMPBSA energies for each trajectory. A total of 100 snapshots are collected from each trajectory over the 8ns production MD runs[29]. For training models on MD trajectory samples, the authors predict the atomic adaptability, which aims to quantify local fluctuations over the trajectory. The authors don’t explore the development of an MMGBSA surrogate model, examine the ranking performance of such a model, or compare the scoring cost versus physics-based methods. The authors do demonstrate an improvement on predicting experimental binding affinities using their MD and QM-derived features[29]. However, this work is limited by the architecture choice of the GNN based on 3D coordinates, likely decreasing sample efficiency and representation robustness[16]. Recently, Liu et al. [31] address the gap of MMPBSA model development using MISATO[29]. The authors frame predicting MMPBSA energetics as a pretraining task, in which the training objective encourages the model to make predictions of the trajectory ensemble-averaged MMPBSA score using ANY frame from that simulation. This objective used by Xu et al. [4] effectively encourages the model to make predictions that are invariant to the subtle changes in the atomic coordinates that vary from snapshots or the docking pose. The authors then consider finetuning their model on the LBA and LEP tasks from ATOM3D[36], where an improvement in performance is achieved by their model on both tasks. While this work represents the first example of “energy-guided pretraining” in the literature, this step is not studied in detail, and it is unclear how well their method generalizes across changes in the 3D structures such as distinguishing multiple poses of the same co-complex structure.
In this Article, we present Surrogate mmGBSA (SurGBSA) for binding 3D structure ranking. Our SurGBSA method comprises (1) the first study of surrogate MMGBSA (or MMPBSA) models trained on snapshot level energetics with an order of magnitude higher snapshot resolution than previously studied and (2) demonstrate a significant improvement in learning the MMGBSA energetics due to physics-informed pretraining. We give the first practical evaluation of MMGBSA surrogate models based on docking pose ranking performance, compared to both the ensemble-averaged MMGBSA and single-point MMGBSA baseline methods. Additionally, we (3) investigate a range of molecular graph encoder modules (GNN, EGNN, EGMN) and their learned representations in detail. Our best-performing SurGBSA model achieves an approximate 6,497 speedup in pose re-ranking while preserving comparable ranking performance to the most efficient MMGBSA baseline, with the difference in correct pose identification being a single sample (). This dramatic efficiency gain enables the high-throughput generation of physics-informed pose ranking, providing the computational foundation necessary for training large-scale MD models that could accelerate virtual screening and drug discovery pipelines. Additionally, in contrast to ensemble-averaged MMGBSA based upon MD, SurGBSA can perform inference without the need to perform additional simulation, greatly reducing the computational resources required for virtual screening. SurGBSA can thus allow for rapid pose ranking comparable to single-point MMGBSA with a dramatic increase in computational efficiency while learning a robust molecular representation. Lastly, we make our implementation and data publicly accessible to support future work in MD foundation model development.
2 Results
2.1 Learning a free-energy surrogate model from MD
We give the schematic of our SurGBSA approach in Figure 1 and give model training details in Section 4.7. We consider each frame of simulation as a unique graph instance paired with a corresponding MMGBSA energy. We train a molecular encoder network with layers to map the 3D atomic structure graph , using atom types and coordinates as inputs, to a set of atomic embeddings which are then average-pooled on the atomic dimension producing a graph summary vector, . The vector is then given as input to a multilayer perceptron (MLP) network which projects to a scalar used to predict the MMGBSA energy.
| Method | Pearson | Spearman |
| GNN | .266 (.065) | .253 (.062) |
| EGNN | .311 (.056) | .293 (.049) |
| EGMN (Scratch) | .436 (.048) | .406 (.044) |
| EGMN (Pretrain) | .573 (.082) | .547 (.078) |
| EGMN (Best) | .702 (.045) | .680 (.043) |
In Table 1 we give the results for the SurGBSA model trained with each molecular encoder network. For fair comparison, we constrained each model to the same training PDB structure sets and use all available MD trajectories corresponding to the top 5 docking poses and the crystal structure. Our results show that the EGMN network, pretrained on molecular dynamics data [27], achieves the best overall performance, by a significant margin. Based on this observation, we extended the training of this model for an additional 500 epochs. Extended training significantly improves both the Pearson and Spearman correlations by 22.5% and 24.5% respectively. To account for possible information leakage of similar structures between each training and test set (5 in total), we computed the protein sequence identity between the PDB structures in each pair of splits (SI Figures 5-9. While we observed a few instances of high similarity, the vast majority of training-testing pairs indicate lower sequence similarity. The high similarity in this context is also not accounting for the 3D conformations of the structure pairs. Additionally, we considered the Tanimoto similarity of the ligands for each PDB structure pair in the respective training/test splits (SI Figures 10-14). Our findings indicate persistent low similarity with only a small number of pairs with high similarity. ProtMD[27] is a model trained to predict MD trajectory ordering and molecular movement between two time steps. The model is evaluated as an additional pre-training step and we considered the possibility of data leakage from the ProtMD pretraining step (SI Figures 15-19). A low overall protein sequence similarity is observed between each of our test sets and the ProtMD pretraining set provided by the authors[27]. This affirms the limited structure redundancy in the MD dataset, by leveraging the structure-similarity analysis used to produce the CASF set of PDBBind [35], which is effective in representative sub-sampling of the available high-quality structures.
In Figure 2, we evaluated our best SurGBSA model and the single-point MMGBSA baseline (SP-GBSA) for accurate docking pose ranking using 10 randomly selected PDB structures. We use the ensemble-average MMGBSA as the ground truth for both methods. We note a high degree of agreement for each set of the trend lines between the SP-GBSA baseline and our SurGBSA model. In several cases, 3/10 for Pearson and 5/10 for Spearman, the ranking produced by SurGBSA improves upon the ranking according to SP-GBSA. Further, when considering the metrics averaged over the PDB structures, SurGBSA shows a slight improvement over the performance of the SP-GBSA baseline which was used as the training target. We believe that this may be attributed to the model effectively learning generalizable features across heterogeneous PDB structures and their conformational dynamics coupled with training on a large collection of MMGBSA energetics that may enable accurate interpolation for test structures supported by a robust representation.
In order to probe the learned representations produced by each molecular graph encoder module, we used Principal Components Analysis (PCA). The intuition for using this approach is that we assume that the model has, over the course of the pre-training and MMGBSA training steps, performed much of the work in learning a non-linear projection of the molecular graphs to their graph embeddings in a lower dimensional space(). By applying linear dimensionality reduction to the learned embeddings, we can assess whether co-complexes naturally separate into interpretable groups, evidence of a robust representation. Figure 3 displays the cumulative distributions of the proportion of variance captured by each principal component. We sought to understand, for each molecular graph encoder module, whether or not training using only crystal structures versus training using MD data (top 5 docking poses and crystal) has an impact on the distribution of explained variance. For example, a representation space for which each orthogonal PCA projection explains exactly the same variance would be an indication of a robust representation. For each test set, we chose the set of the static structures for the top 5 docking poses (i.e. frame 0) for each PDB structure, extracted their graph embeddings , then performed PCA and collected the explained variance ratios. We averaged the ratios across all 15 of our independent trials for each model hyperparameter configuration (5 folds, 3 seeds). Each distribution sums approximately to 1 with infinitesimal error. We observe in the case of training using only crystal structures and the corresponding SP-GBSA energy, nearly all of the encoder modules we considered contain most of the explained variance in a single dimension, with the exception of our best EGMN encoder which is slightly below the other models for the first component. When considering the models trained on MD trajectories, we note a stark difference in the shapes of the curves that indicates that the explained variance is more uniformly spread throughout the dimensions, although the models we trained from scratch still exhibit a large proportion of the variance explained by the first few components. The order of the curves precisely follows the dimension of model complexity, ranging from the most simple GNN to the most complex physics-informed EGMN encoders. Our extended training of the pre-trained EGMN shows the best result, with the first component responsible for only 30% of the variance and with the remaining variance more uniformly spread relative to all other encoder modules. We visualized the PCA components for each encoder module in SI Figures 20-44.
| Model | Spearman | ||
| GNN | .17 | -68.5 | 1.0 |
| EGNN | .27 | -50.0 | 58.8 |
| EGMN (scratch) | .30 | -44.4 | 76.5 |
| EGMN (pretrain) | .33 | -38.9 | 94.1 |
| EGMN (best) | .45 | -16.7 | 164.7 |
| SP-GBSA | .54 | 1.0 | 217.6 |
| Model | Correct Top Pose % |
| GNN | 30.0 |
| EGNN | 33.1 |
| EGMN (scratch) | 38.0 |
| EGMN (pretrain) | 40.1 |
| EGMN (best) | 44.7 |
| SP-GBSA | 45.1 |
In Figure 4 and Table 2 we detail the comparison in docking pose re-ranking with respect to the MMGBSA score based on all docking pose MD snapshots versus SurGBSA, using the various molecular encoder networks considered in this work, and the SP-GBSA baseline. Although our results show that SP-GBSA achieves the highest mean Spearman taken over all PDB structures in each test set, the difference in this ranking performance versus our best SurGBSA model is . Compared to using our baseline GNN encoder module, the improvement by our best SurGBSA model is significant at . Lastly, in Table 3 we give the results for correct pose identification. For each PDB structure, we presented each model with the five corresponding docking poses we generated. The poses are then ranked according to both the model prediction and to their ensemble-averaged MMGBSA energies. If the model top ranking agrees with the MMGBSA ranking, then we count this as a success. We again note a trend in terms of model complexity with our best overall method using the physics-informed EGMN extended training run. When comparing to the baseline re-scoing method, SP-GBSA, the difference is only a single test sample.
2.2 Efficiency of SurGBSA
| Method | Host | Scoring (s) | Speedup |
| MMGBSA [37] | CPU (Dane) | 11.079 | 0.2 |
| SP-GBSA [37] | CPU (Dane) | 1.949 | 1.0 |
| SurGBSA | AMD MI300A (Tuolumne) | 0.002 | 974.5 |
| SurGBSA | Nvidia H100 (Matrix) | 0.0003 | 6496.7 |
The MMGBSA score itself is relatively expensive compared to docking scoring functions as it requires MD simulation to sample conformations to improve its accuracy [26]. In Section 2.1 we discussed the ability to accurately recapitulate the MMGBSA score through a physics-informed surrogate model. Next, the surrogate model’s computational efficiency is compared to the MMGBSA score. In Table 4 we provide performance analysis collected on HPC clusters (Dane, Matrix, and Tuolumne). We considered two versions of MMPBSA.py [37] included in the Amber 22 MD simulation toolkit, a single core implementation and a multithreaded MPI implementation (MMPBSA.MPI.py). We only considered the Generalized Born (GB) solvation model as it is more efficient and serves as a more competitive baseline while scaling reasonably to the full dataset we consider in this work. As the MMGBSA calculation is parallelized over the frame dimension, SP-GBSA can only take advantage of a single core of the CPU per frame. Our results show that for the latency required to score a single docking pose versus the most competitive baseline, SP-GBSA, our SurGBSA model can achieve a speed-up improvement of approximately 6,497. Compared to running the parallel MPI version of the MMGBSA scoring function, using all snapshots from an MD trajectory, the speedup is approximately 36,930. Given the improved efficiency in conjunction with docking pose ranking performance as compared to the SP-GBSA and MMGBSA baselines, we believe that SurGBSA is capable of serving as a scalable method for docking pose re-scoring that can be used to supplement more expensive methods, including MMGBSA based on MD, by providing an efficient and comparatively accurate predictor. SurGBSA does not require MD for inference; thus, it can be applied to significantly larger datasets of docking poses by several orders of magnitude, improving the size of the tractable search space. This scalability allows SurGBSA to augment the more expensive MD-based scoring methods as a pre-processing step for ranking 3D structures of protein-ligand co-complexes.
3 Conclusion
SurGBSA is a new surrogate model for physics-based scoring of binding affinity of protein-ligand co-complex structures. Extending previous physics-informed pre-training, the new model learns an accurate and efficient mapping between molecular structure phase space and the corresponding binding free energy surface using the MMGBSA score. We validate our approach using a rigorous train/test split of representative protein-ligand structures and cross-validation with multiple random seeds for each evaluated model architecture. Results demonstrate that the SurGBSA model, which combines ProtMD[27] physics-informed pre-training with learning a mapping between molecular conformation and binding free energy produces a more robust representation compared to methods not trained on molecular dynamics trajectories. The SurGBSA model differs from the baseline physics-based single point (SP) GBSA calculation by a single ranking (i.e. ), when evaluated on a challenging pose ranking problem, with approximately a 6,497 speedup.
Future work will investigate improving the pre-training stage using recently published datasets such as ATLAS [38] which contains 1390 standardized MD simulations for a variety of PDB structures. Improving the inductive biases of the EGMN architecture to more efficiently model the molecular structure dynamics could improve scaling to larger structures. Adding new pre-training tasks and more training examples should continue to improve SurGBSA model accuracy in downstream applications. Physics-informed pretraining can contribute to improving the collection of predictive protein-ligand interaction models.
4 Methods
In this section we first describe the MD dataset used as the basis for the SurGBSA model development. We then introduce background on the MMGBSA method and describe the tools used to compute the MMGBSA data used in this work. We then provide descriptions of each graph encoder network architecture and the MLP architecture used for learning the MMGBSA score based on the graph encoder output embeddings. Finally, we describe the model hyperparameter optimization and training.
4.1 Dataset
We use the publicly available collection of protein-ligand co-complexes available in the PDBBind 2019. In order to collect a representative subsampling of the full PDBBind, which contains over 18k unique co-complex structures, we use the PDBBind Core set which is a subset of 285 structures that span the structural space present in the refined set. We successfully computed 5 docking poses for 244 structures out of the full 285, resulting in 6 structures for each of the 244 co-complexes. Thus we have 1464 total “poses” including the crystal structure pose. We used Amber to collect 10 nanosecond atomistic molecular dynamics simulations for each of the 1464 poses with snapshots collected every 10 picoseconds, resulting in 1000 frames of simulation per each of the 1464 poses. For each of the snapshots, we collected an MMGBSA score to estimate the binding free energy . Thus, in total, our full MD dataset thus contains 1,464,000 frames of simulation with corresponding MMGBSA scores for each frame.
To prepare the dataset for learning, we filter the simulation frames based upon the pocket-ligand co-complex structure to the starting frame, excluding all frames beyond to account for instabilities in the simulations while retaining meaningful fluctuations and conformational changes. We also filter out frames with MMGBSA scores greater than or equal to -10 and with MMGBSA scores less than -100. We then scaled the MMGBSA scores by dividing by a constant factor of -10, as we found that it leads to better numerical stability during training. We only keep the atoms within a radius of 6 angstroms from the ligand and use a maximum of 600 atoms while we keep the full ligand in each frame. We used the MDAnalysis [39, 40] python library to process the MD simulations. We aligned each trajectory with the respective crystal structure simulation and collected the RMSD for each frame with respect to the crystal pose. We filter out all water molecules, metals, and ions. To measure generalization performance, we considered a K-fold split of the core set complexes with , the default value of the scikit-learn utility, resulting in approximately 80% of the data used for training and 20% for testing. In total, after filtering based on the MMGBSA score and pocket-ligand co-complex starting point RMSD, we retain a total of 1,070,272 frames of simulation out of the 1.4 million collected from the MD simulations (). We encode the atom types and positions as input for the models.
4.2 MMGBSA
The MMGBSA method is an end-point free energy method that is more computationally efficient compared to Alchemical free energy (AFE) methods while being more accurate than docking scoring functions [26]. Due to these properties, MMGBSA has found application for evaluating docking poses and predicting binding affinities [26]. For our baselines, we use the MMPBSA.py script provided in the AmberTools22 software package [41]. We calculated single-point MMGBSA scores for each simulation frame in our dataset. We formed two distinct MMGBSA baselines; (1) SP-GBSA: uses the docked or crystal pose and the corresponding single-point MMGBSA score, and (2) MMGBSA which is based on the average of the single-point MMGBSA scores for each docking pose or crystal structure MD trajectory. SP-GBSA serves to function as a realistic efficient approximation to conserve computational resources when selecting ligands and their poses for further MD simulation. MMGBSA serves to represent the most accurate version of the re-scoring function based on all available conformational dynamics.
4.3 Graph Neural Networks
Graph Neural Networks [42, 43, 44] operate on undirected structured graph data and are a natural choice for molecular representation learning [42, 45, 20]. We based our GNN on the implementation provided by Satorras et al. [16]. Given a graph and node embeddings where , the Graph Convolution Layer (GCL) computes a permutation equivariant transformation on the node embeddings . The GCL is defined by the following equations:
4.4 E(n) Equivariant Graph Neural Networks
The Equivariant Graph Convolutional Layer (EGCL) [16] improves the GCL by introducing equivariant operations, which improves sample efficiency when processing different conformations of the same molecule(s) as compared to the GNN. We based our EGCL on the implementation provided by Satorras et al. [16]. Given the set of node embeddings , edge information , coordinate embeddings where , the EGCL layer computes a transformation on the node embeddings and coordinates . The EGCL is defined by the following equations:
where is a scaling constant, is a learned differentiable function, and is the relative squared distance between and . From the output of the EGCL layer, the node embeddings are invariant and the node coordinates are equivariant to . The EGNN improves sample efficiency in contrast to the GNN in that the equivariant properties allow for the model to recognize different conformations of the same molecule, thus improving generalization under rotation and translations.
4.5 Equivariant Graph Matching Networks
The Equivariant Graph Matching Network (EGMN) architecture implemented by Wu et al. [27] is closely related to the Independent E(3)-Equivariant Graph Matching Network (IEGMNs) [46]. Wu et al. [27] treat the time dimension in MD simulations by considering each time-dependent conformation as distinct molecular graphs. Specifically, the EGMN assumes distinct graphs , and additional cross-graph edges . The EGMN takes as input the set of node embeddings and coordinates at time and layer then computes the transformation , defined by the following equations:
where corresponds to L2 normalized euclidean distances between nodes and in a given graph . , , , and are learnable parameterized functions. Specifically, if otherwise when , where and are learnable parameterized functions specific to inter-graph or intra-graph edge types. The attention coefficient between the nodes and is calculated as:
where and are trainable neural networks.
4.6 Supervised models
To learn the MMGBSA score, we first aggregated the learned embeddings extracted from each graph encoder network:
where is the number of graph convolution layers and , are the initial node embeddings and coordinates, respectively. In the case of the GNN encoder, we treat as it does not output a coordinate transformation. We discard and keep only the node embeddings for the rest of the computation. To compute the embedding of the graph , we use average pooling over the atomic dimension:
where . Then, we use an MLP network to map to the scalar output:
where and are weight matrices, and are bias vectors, and is the ReLU non-linearity.
4.7 Model training and optimization
We considered a grid search over various settings of the hyperparameters. We considered batch sizes of 64, 128, and 256. We considered learning rates of and . We train all models using the AdamW optimizer with no weight decay in order to constrain the hyperparameter optimization and focus our study on the learning dynamics induced by the choice of learning rate and batch size instead. We considered 5 distinct random splits of the CASF-2016 complexes, which represent a representative sub-sampling of the PDBBind 2019 refined set structures. We trained a model, independently, for each unique configuration of hyperparameters on each of the 5 folds. Further, we replicated each experiment using 3 random seeds, thus for each configuration of hyperparameters we have a total of 15 independent training runs. In total, we conducted 5400 training runs for our study using the PyTorch [47] Distributed Data-Parallel (DDP) backend (v. 2.5.1), AMD ROCm 6.2, on the Tuloumne HPC cluster equipped with the AMD MI300A Accelerated Processing Unit (APU)[48]. In contrast to the GPU accelerator where the CPU and GPU access disjoint memory spaces, the APU combines on-chip memory resources that are shared across both the CPU and GPU, improving memory bandwidth between the processing units[48]. For training runs using only static information (i.e. crystal structure and docking poses) we used 1 node with 4 total APUs per job. For training runs using the MD conformations, we used 4 nodes per job with a total of 16 APUs per job. We trained all models for a maximum of 100 epochs, regardless of the dataset. Using the best model configuration after 100 epochs, the ProtMD pre-trained EGMN encoder network, we extended training to 600 epochs and report this combination as our best SurGBSA model. We treat the graphs from different frames independently, i.e. as separate graphs and train all layers of the constituent molecular graph encoder and MMGBSA prediction head using the MSE loss. We report the results using the models with the hyperparameters that achieve the best average spearman correlation, averaged by PDB structure, on the test set.
Acknowledgments
This research was supported by funds from the UC National Laboratory Fees Research Program of the University of California, Grant Number L23GF6259. This work was supported in part by CRISP and PRISM, centers in JUMP 1.0 and 2.0, SRC programs sponsored by DARPA, SRC #236160. Computing support for this work came from the Lawrence Livermore National Laboratory (LLNL) Institutional Computing Grand Challenge program. This work was funded in part by the Defense Threat Reduction Agency (DTRA) HDTRA1036045. Part of this research was also supported by the American Heart Association under CRADA TC02274-4. All work performed at Lawrence Livermore National Laboratory is performed under the auspices of the U.S. Department of Energy under Contract DE-AC52-07NA27344, LLNL-JRNL-2010833-DRAFT.
[…]
Supplementary information
5 Dataset Splits
| Fold | Train | Val | Test |
| 0 | 172 | 20 | 49 |
| 1 | 173 | 20 | 48 |
| 2 | 173 | 20 | 48 |
| 3 | 173 | 20 | 48 |
| 4 | 173 | 20 | 48 |
6 Model Architecture Details
| Encoder | # Params. | dim | |
| GNN | 445,312 | 128 | 4 |
| EGNN | 495,872 | 128 | 4 |
| EGMN | 5,237,445 | 128 | 6 |
| MLP | 8,321 | - | - |
7 Sequence similarity of each training set versus the corresponding test set by fold
8 Ligand Similarity Clustermaps
9 Sequence similarity of the ProtMD training set versus each test set
10 PCA Projections of Model Representations